# LEAD: Towards Learning-Based Equity-Aware Decarbonization in Ridesharing Platforms

### v1

@@ Abstract
Ridesharing platforms such as Uber, Lyft, and DiDi have grown in popularity due to their on-demand availability, ease of use, and commute cost reductions, among other benefits. However, not all ridesharing promises have panned out. Recent studies demonstrate that the expected drop in traffic congestion and reduction in greenhouse gas (GHG) emissions have not materialized. This is primarily due to the substantial distances traveled by the ridesharing vehicles without passengers between rides, known as deadhead miles. Recent work has focused on reducing the impact of deadhead miles while considering additional metrics such as rider waiting time, GHG emissions from deadhead miles, or driver earnings. Unfortunately, prior studies consider these environmental and equity-based metrics individually despite them being interrelated.
In this paper, we propose a Learning-based Equity-Aware Decarabonization approach, LEAD, for ridesharing platforms. LEAD targets minimizing emissions while ensuring that the driver’s utility, defined as the difference between the trip distance and the deadhead miles, is fairly distributed. LEAD uses reinforcement learning to match riders to drivers based on the expected future utility of drivers and the expected carbon emissions of the platform without increasing the rider waiting times. Extensive experiments based on a real-world ride-sharing dataset show that LEAD improves fairness by 2$\times$ when compared to emission-aware ride-assignment and reduces emissions by 70% while ensuring fairness within 66% of the fair baseline. It also reduces the rider wait time, by at least 40%, compared to various baselines. Additionally, LEAD corrects the imbalance in previous emission-aware ride assignment algorithms that overassigned rides to low-emission vehicles.
1 Introduction
@@
Road transportation significantly contributes to global energy consumption and greenhouse gas (GHG) emissions [8]. Worldwide, the transportation sector is the fourth-largest source of GHG emissions and is the largest in the United States [42, 29]. In 2021, transportation accounted for 37% of global CO2 emissions [17]. That same year, the United States emitted approximately 5 billion metric tons of carbon dioxide, accounting for about 13.49% of global emissions, surpassing the combined emissions of all 28 European Union countries [7]. Due to these environmental impacts, there is an effort towards exploring transportation systems that transport people in an eco-friendly manner [33].
@@
In recent years, shared mobility options have emerged as an attractive transport solution, thanks to the widespread integration of smartphones into daily life. These services enable users to access transportation on demand without owning a vehicle [34, 4, 26, 27]. As a result, ridesharing is seen as a sustainable transportation option that can help reduce emissions and traffic congestion. Unfortunately, these expectations have not materialized and ridesharing services have been linked to increased traffic congestion, higher emissions, distracted driving, and negative social equity implications [14, 30, 3]. The increase in traffic congestion and emissions results from two key by-products of ridesharing. First, ridesharing availability can reduce the use of public transportation. Second, ridesharing vehicles travel a significant fraction of their total miles without a passenger on their way to pick up a passenger or come back after dropping one [23, 36, 31]. These non-passenger miles are referred to as deadhead miles. Previous studies show that deadhead miles can account for 19% to 41% of total miles while increasing emissions by up to 90% as compared to the personal vehicle use [15, 46].
@@
There has been significant recent work on optimizing the deadhead miles in ridesharing ecosystems [18, 32] while considering additional rider-specific metrics such as the wait time [47, 19] or platform-wide objectives such as GHG emissions [32, 49]. Other studies aim to enhance overall rider satisfaction by minimizing travel costs and waiting times [6, 20, 37]. However, these studies often ignore fairness from the drivers’ perspective, which can exacerbate social inequities. The studies from a diver’s perspective typically focus on enhancing driver satisfaction and operational efficiency by optimizing routes and maximizing a given definition of utility [21, 40]. While some of those studies also target fairness in earnings across drivers [35], they ignore the rider’s wait time and system-wide emissions. As a result, the ride-assignment algorithms proposed by prior cannot be used in practice to promote sustainable and equitable growth in ridesharing platforms.
@@
In this work, we consider the problem of holistic ride assignment that balances the objectives of the system (low emissions), the driver (fairness in earnings), and the rider (low waiting time). The goal is to minimize system-wide emissions, including emissions from ride and deadhead miles, while maximizing fairness in total utility across all trips between drivers without impacting the rider’s waiting time. The utility of a trip for a driver is defined as the trip distance minus the deadhead distance. Given this setup, [32] and [35] represent our most relevant related works. In [32], the authors propose an online threshold-based ride assignment algorithm (TORA) that targets minimizing the system-wide emissions while optimizing the waiting time of the riders. TORA achieves a highly favorable tradeoff between emissions (up to 60% reduction) and rider’s wait time (4% increase). However, to reduce emissions, TORA assigns a disproportionate fraction of rides to low-emission vehicles, resulting in high unfairness among drivers. For example, it may assign up to 65% of the rides to EVs despite them constituting only 25% of the fleet. In [35], the authors propose LAF that targets maximizing utility, defined as the total revenue generated by the system while minimizing unfairness in the driver’s equitable earnings. While taking a driver-centric approach, LAF completely ignores system-wide emissions and the rider’s wait time.
@@
In this paper, we propose the Learning-Based Equity-Aware Decarbonization (LEAD) algorithm, which aims to minimize total system-wide emissions and maximize fairness in drivers’ accumulated utilities across trips. To reduce system-wide emissions, LEAD must consider not only the drivers that are available at present but also the ones that will become available in the near future. To do so, LEAD explicitly accounts for the dependencies among ride assignments by using reinforcement learning to develop future-aware ride assignment strategies. Using a real-world ridesharing dataset, we analyze how LEAD and other representative baselines [32, 35] perform using metrics, such as the reduction in emissions, the quality of service for riders measured as wait time, and fairness for drivers, and how gracefully they navigate the trade-offs between these metrics. Our contributions can be summarized as follows:
@@
  • We implemented LEAD on a simulation testbed and evaluated its performance using real data from RideAustin [28], a nonprofit ridesharing service. In addition, we compare the performance of LEAD with heuristic and state-of-the-art algorithms used in [32, 35]. Our experimental results show that, in practical scenarios, our algorithm not only significantly improves the fairness of utility distribution among drivers but also substantially reduces emissions, compared to the current state-of-the-art methods.
2 Related Work
@@
Ridesharing Optimization Approaches. The development of ridesharing systems has introduced unique challenges, with the primary focus being the assignment of ride requests to drivers. Some challenges in this area include improving vehicle utilization across multiple ride requests and demands [48, 25, 22], developing algorithms for route planning [10, 43], and analyzing and integrating temporal and spatial patterns to predict the arrival time of riders’ requests [16, 11]. Some other works in this domain also consider different optimization objectives such as maximizing drivers’ profit [45, 2, 51], while others aim to minimize travel costs [41, 44] or reduce riders’ waiting times [50, 1, 47].
@@
In [18] , authors attempt to reduce trip-level deadhead miles by leveraging hour-ahead trip demand predictions and using a heuristic approach to driver assignment. However, this focus on reducing trip-level deadhead miles does not always result in a reduction in system-wide deadhead miles and emissions, as these outcomes also depend on factors such as the fuel efficiency of vehicles and traffic conditions. Another work  [32] focuses on minimizing emissions from deadhead miles while also reducing rider waiting times, ensuring no degradation in user satisfaction. However, a ride assignment system focused only on reducing emissions can negatively impact the quality of service (QoS) for riders and drivers. For instance, prioritizing emission reduction as the primary objective can lead to a preference for electric or low-emission vehicles over high-emission ones. This bias is unfair to drivers with high-emission vehicles, who often belong to lower-income communities and may face reduced earning opportunities. Additionally, an emissions-aware ride assignment may allocate trips with inherently longer deadhead miles to low-emission vehicles. This increases the deadhead-to-trip ratio, decreasing these vehicles’ overall efficiency and service quality.
@@
Another key challenge is the lack of consideration for the temporal dependencies between current and future assignments. This oversight can impact the optimization of utility and fairness. Many algorithms in this area attempt to provide theoretical performance guarantees, but they typically rely on myopic assumptions [47]. For instance, they frequently neglect the dependency of assignments on past and future decisions, which can lead to sub-optimal assignments. This shortsightedness fails to account for the dynamic and evolving nature of ridesharing systems, where current ride assignments can significantly impact future states and the system’s overall efficiency. An effective alternative approach involves leveraging historical data to optimize ride assignments alongside integrating reasonable assumptions.
@@
Fairness in Ridesharing. Fairness is a critical factor in matching scenarios, particularly in multi-sided markets where preferences play a significant role [12, 24, 9]. In the context of ridesharing, recent literature has raised concerns about the fairness of the assignments. Brown et al. [5] shows the unfair treatment of riders by ridesharing companies, which leads to higher trip cancellation rates for a few groups of riders. Another study indicates that income inequality among ridesharing drivers may prevent them from earning a living wage [13]. Another work aims to optimize long-term efficiency and fairness in ridesharing platforms through joint order dispatching and driver repositioning [38]. Another recent work [32] showed a trade-off between reducing carbon emissions and sacrificing the fairness of utilities among drivers. Similarly, another study [35] demonstrated optimization in driver utilities and fairness without considering emissions. This study used reinforcement learning techniques to evaluate the expected achievable utilities in different city areas. Despite the success of this approach in ensuring equity among drivers’ utilities, it did not address eco-friendly ride assignments with low carbon emissions.
@@
Ensuring fair assignments in ridesharing systems is challenging due to drivers’ and riders’ inherent spatiotemporal dynamics. Drivers can earn varying amounts of utility based on their assignments, while riders may experience differing waiting times. These disparities can lead to perceptions of unfairness, complicating the assignment process. To tackle these challenges, our work introduces a new fairness metric for driver earnings that considers the spatiotemporal dynamics inherent in ridesharing systems. This metric is designed to ensure a more equitable distribution of earnings among drivers, considering the dynamic nature of the system. Building on this metric, we have developed a reinforcement learning-based approach assignment algorithm that reduces carbon emissions while providing a fair distribution of driver utilities compared to the state-of-the-art assignment algorithm [32]. The algorithm leverages historical data and dynamic learning to create effective and equitable assignment strategies, addressing immediate and long-term objectives in the ridesharing ecosystem.
3 Problem Formulation
@@
In ridesharing platforms, riders set the pickup and dropoff locations for a ride request; then, the platform matches the request to one of the available drivers and estimates the trip duration and waiting time based on factors such as trip distance, traffic congestion, and the distance between the assigned driver and the rider’s pickup location. An available driver in the system refers to a driver who is either idle or soon to be available within the specified time to serve a request. The emission associated with a ride request $r$ is influenced by the unit emissions of the driver’s vehicle, the deadhead distance, the trip distance, and additional factors such as traffic congestion, road conditions, etc. For simplicity, we estimate the carbon emissions produced during the servicing of the ride request $r$ by driver $v$ using the equation below.
$E_{v,r}=(d_{v,r}^{(D)}+d_{r}^{(T)})\cdot e_{v},$ (1)
@@
$U_{v,r}=\sum_{r^{\prime}\in\mathcal{V}_{r}}(d_{r^{\prime}}^{(T)}-d_{v,r^{\prime}}^{(D)})\cdot x_{v,r^{\prime}},$ (2)
@@
Table 1: Summary of important notations.
Notation Description
$\mathcal{V}$ Set of all drivers
$\mathcal{R}$ Set of all ride requests
$\mathcal{V}_{b}$ Set of available drivers during batch $b$
$\mathcal{R}_{b}$ Set of ride requests in batch $b$
$p_{r}$ Pickup location of ride request $r$
$q_{r}$ Dropoff location of ride request $r$
$c_{r}$ Created time of ride request $r$
$a_{r,v}$ Dropoff time of ride request $r$ when served by driver $v$
$E_{v,r}$ Emission of serving request $r$ by driver $v$
$U_{v,r}$ Utility history of driver $v$ before serving ride request $r$
$d^{(T)}_{r}$ Trip distance of ride request $r$
$d^{(D)}_{v,r}$ Deadhead distance of serving ride request $r$ by driver $v$
$d_{t,D}$ Deadhead distance of $t^{th}$ trip served by the agent
$d_{t,T}$ Trip distance of $t^{th}$ trip served by the agent
$V_{D}(s)$ Value function at state $s$ associated to deadhead distances
$V_{T}(s)$ Value function at state $s$ associated to trip distances
$e_{v}$ Produced carbon emission per mile of the vehicle of driver $v$
@@
$\text{Fairness}:=-\max_{v}(U_{v})$
$\quad\ \ \ \ +\min_{v}(U_{v}),$
s.t., $\sum_{v}x_{v,r}\in\{0,1\}\qquad\forall r,$ (3e)
@@
vars., $\ \ \quad x_{v,r}\in\{0,1\}\qquad\forall v,r.$ (3f)

### v2

@@ Abstract.
Ridesharing platforms such as Uber, Lyft, and DiDi have grown in popularity due to their on-demand availability, ease of use, and commute cost reductions, among other benefits. However, not all ridesharing promises have panned out. Recent studies demonstrate that the expected drop in traffic congestion and reduction in greenhouse gas (GHG) emissions have not materialized. This is primarily due to the substantial distances traveled by the ridesharing vehicles without passengers between rides, known as deadhead miles. Recent work has focused on reducing the impact of deadhead miles while considering additional metrics such as rider waiting time, GHG emissions from deadhead miles, or driver earnings. However, most prior studies consider these environmental and equity-based metrics individually despite them being interrelated. In this paper, we propose a Learning-based Equity-Aware Decarabonization approach, LEAD, for ridesharing platforms. LEAD targets minimizing emissions while ensuring that the driver’s utility, defined as the difference between the trip distance and the deadhead miles, is fairly distributed. LEAD uses reinforcement learning to match riders with drivers based on the expected future utility of drivers and the expected carbon emissions of the platform without increasing the rider waiting times. Extensive experiments based on a real-world ridesharing dataset show that LEAD improves the defined notion of fairness by 150% when compared to emission-aware ride-assignment and reduces emissions by 14.6% while ensuring fairness within 28–52% of the fairness-focused baseline. It also reduces the rider wait time, by at least 32.1%, compared to a fairness-focused baseline.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 23-26, 2025; Athens, Greeceisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Road transportation significantly contributes to global energy consumption and greenhouse gas (GHG) emissions (Davis and Boundy, 2021). Worldwide, the transportation sector is the fourth-largest source of GHG emissions and is the largest in the United States (U.S. Environmental Protection Agency, 2024; Sadati and Çatay, 2021). In 2021, transportation accounted for 37% of global CO2 emissions (International Energy Agency, 2022). That same year, the United States emitted approximately 5 billion metric tons of carbon dioxide, accounting for about 13.49% of global emissions, exceeding the combined emissions of the 28 countries of the European Union (Climate.gov, 2024). Due to these environmental impacts, there is an effort to explore transport systems that move people in an eco-friendly manner (Schnieder, 2023).
@@
In recent years, shared mobility options have emerged as an attractive transport solution due to the widespread integration of smartphones into daily life. These services enable users to access transportation on demand without owning a vehicle (Shaheen et al., 2016; Bongiovanni et al., 2022; Ma et al., 2020, 2022). As a result, ridesharing is seen as a sustainable transportation option that can help reduce emissions and traffic congestion. Unfortunately, these expectations have not materialized and ridesharing services have been linked to increased traffic congestion, higher emissions, and negative social equity implications (Henao, 2017; Safaeian et al., 2023; Barrios et al., 2023). The increase in traffic congestion and emissions results from two key by-products of ridesharing. First, ridesharing availability can reduce the use of public transportation. Second, ridesharing vehicles travel a significant fraction of their total miles without a passenger on their way to pick up a passenger or return after dropping one (Loa et al., 2023; Shi et al., 2021a; Sahebdel et al., 2023). These non-passenger miles are referred to as deadhead miles. Previous studies show that deadhead miles can account for 19% to 41% of total miles while increasing emissions by up to 90% compared to personal vehicle use (Henao and Marshall, 2019; Wenzel et al., 2019).
@@
Recent work has focused on optimizing deadhead miles in ridesharing ecosystems (Kontou et al., 2020; Sahebdel et al., 2024) while considering additional rider-specific metrics such as wait time (Xu et al., 2019; Lee et al., 2004) or platform-wide objectives such as GHG emissions (Sahebdel et al., 2024; Yan et al., 2020). Other studies aim to improve overall rider satisfaction by minimizing travel costs and waiting times (Cheng et al., 2017; Levinger et al., 2020; Singh et al., 2021), but they often ignore fairness from the drivers’ perspective, which can exacerbate social inequities. The studies from a driver’s perspective typically focus on enhancing driver satisfaction and operational efficiency by optimizing routes and maximizing a given definition of utility (Li et al., 2022; Tang et al., 2019). While some of those studies target fairness in earnings across drivers (Shi et al., 2021b), they ignore the rider’s waiting time and system-wide emissions. As a result, the ride-assignment algorithms proposed by prior work cannot be used in practice to promote sustainable and equitable growth in ridesharing platforms.
@@
In this work, we consider the ride assignment problem that balances the objectives of the system (low emissions), the driver (fairness in earnings), and the rider (low waiting time). The goal is to minimize system-wide emissions, including emissions from ride and deadhead miles, while maximizing fairness in total utility across all trips between drivers without impacting the rider’s waiting time. The utility of a trip for a driver is defined as the trip distance minus the deadhead distance. Given this setup, (Sahebdel et al., 2024) and (Shi et al., 2021b) represent our most relevant related works. In (Sahebdel et al., 2024), the authors propose an online threshold-based ride assignment algorithm (TORA) that targets minimizing the system-wide emissions while optimizing the waiting times. TORA achieves a highly favorable tradeoff between emissions (up to 60% reduction) and rider’s wait time (4% increase). However, to reduce emissions, TORA assigns a disproportionate fraction of rides to low-emission vehicles, resulting in high unfairness among drivers. For example, it may assign up to 65% of the rides to EVs despite them constituting only 25% of the fleet. In (Shi et al., 2021b), the authors propose LAF that targets maximizing utility, defined as the total revenue generated by the system, while minimizing unfairness in the driver’s equitable earnings. While taking a driver-centric approach, LAF completely ignores system-wide emissions and the rider’s wait time.
To achieve the objectives of reducing system-wide emissions and ensuring fairness for drivers, we leverage reinforcement learning (RL). RL’s potential to handle sequential decisions and adapt based on feedback makes it a powerful tool for navigating dynamic and intricate environments. RL provides a framework in which agents learn to optimize their actions within an environment through trial and error, making it applicable to a wide range of application domains (Sales-Ortiz and Ardakanian, 2024; Han et al., 2022; Pigott et al., 2022; Wang et al., 2024). In the setup of RL in the context of ridesharing, an agent performs an action that directly impacts the environment, which could include factors related to the city, drivers, or riders. In response, the agent receives feedback, such as the successful trip completion or the revenue earned, and uses this information to refine its future decision-making. Its ability to capture and understand the intricate spatial and temporal dynamics of demand patterns has shown great promise in solving operational inefficiencies commonly faced in ridesharing systems, such as vehicle dispatching, matching riders with drivers, and optimizing route planning in real time (Cao et al., 2020; Asadi and Pinkley, 2021).
@@
In this paper, we propose the Learning-Based Equity-Aware Decarbonization (LEAD) algorithm, which aims to minimize total system-wide emissions and maximize fairness in drivers’ accumulated utilities across trips. To reduce system-wide emissions, LEAD must consider the drivers available at present and the ones that will become available in the near future. To do so, LEAD explicitly accounts for the dependencies among ride assignments by using reinforcement learning to develop future-aware ride assignment strategies. Using a real-world ridesharing dataset, we analyze how LEAD and other representative baselines (Sahebdel et al., 2024; Shi et al., 2021b) perform using metrics, such as reduction in emissions, quality of service for riders measured as wait time, and fairness for drivers, and how gracefully they navigate the trade-offs between these metrics. Our contributions can be summarized as follows:
@@
  • We implemented LEAD on a simulation testbed and evaluated its performance using real data from RideAustin (RideAustin, 2017), a nonprofit ridesharing service. In addition, we compare the performance of LEAD with heuristic and state-of-the-art algorithms used in (Sahebdel et al., 2024; Shi et al., 2021b). Our experimental results show that, in practical scenarios, our algorithm not only significantly improves the fairness of utility distribution among drivers but also substantially reduces emissions, compared to the current state-of-the-art methods.
2. Related Work
@@
Ridesharing Optimization Approaches. The development of ridesharing systems has introduced unique challenges, with the primary focus being the assignment of ride requests to drivers. Some challenges in this area include improving vehicle utilization across multiple ride requests and demands (Xu et al., 2020; Luo et al., 2019; Li et al., 2019), developing algorithms for route planning (Duan et al., 2020; Wang et al., 2020), and analyzing and integrating temporal and spatial patterns to predict the arrival time of riders’ requests (Hong et al., 2020; Fang et al., 2021). Some other works in this domain also consider different optimization objectives such as maximizing drivers’ profit (Wang et al., 2019; Asghari and Shahabi, 2018; Zhang et al., 2017), while others aim to minimize travel costs (Tong et al., 2016; Wang et al., 2018) or reduce riders’ waiting times (Yengejeh and Smith, 2021; Anthony and Chung, 2014; Xu et al., 2019).
@@
In (Kontou et al., 2020) , authors attempt to reduce trip-level deadhead miles by leveraging hour-ahead trip demand predictions and using a heuristic approach to driver assignment. However, this focus on reducing trip-level deadhead miles does not always result in a reduction in system-wide deadhead miles and emissions, as these outcomes also depend on factors such as the fuel efficiency of vehicles and traffic conditions. Another work  (Sahebdel et al., 2024) focuses on minimizing emissions from deadhead miles while also reducing rider waiting times, ensuring no degradation in user satisfaction. However, a ride assignment system focused only on reducing emissions can negatively impact the quality of service for riders and drivers. For example, a naive algorithm focused primarily on emission reduction may favor electric or low-emission vehicles over high-emission ones, leading to unfair ride assignments for drivers with high-emission vehicles. These drivers often belong to lower-income communities and may face diminished earning opportunities as a result. Additionally, an emissions-aware ride assignment may allocate trips with longer deadhead miles to low-emission vehicles. This increases the deadhead-to-trip ratio, decreasing these vehicles’ overall efficiency and service quality.
@@
Another key challenge is the lack of consideration for the temporal dependencies between current and future assignments. This oversight can impact the optimization of utility and fairness. Many algorithms in this area attempt to provide theoretical performance guarantees, but they typically rely on myopic assumptions (Xu et al., 2019). For instance, they frequently neglect the dependency of assignments on past and future decisions, which can lead to sub-optimal assignments. This shortsightedness fails to account for the dynamic and evolving nature of ridesharing systems, where current ride assignments can significantly impact future states and the system’s overall efficiency. A practical alternative approach involves leveraging historical data to optimize ride assignments alongside integrating reasonable assumptions.
Reinforcement Learning in Ridsharing. Numerous prior studies have highlighted the potential and advantages of utilizing reinforcement learning. Several comprehensive surveys on reinforcement learning (RL) for intelligent transportation systems exist, with most focusing on applications such as traffic signal control, autonomous driving, and routing optimization (Haydari and Yılmaz, 2020; Yau et al., 2017). Song et al. (Song et al., 2020) conduct a case study on ridesharing in Seoul, utilizing a tabular Q-learning agent to optimize spatiotemporal pricing. Their extensive simulations focus on the issue of driver refusals to transport passengers to low-demand areas, which results in longer wait times for riders in these regions. Additionally, they investigate how surge pricing can mitigate the challenges faced by these marginalized zones—areas where taxis are typically scarce—and contribute to improving spatial equity throughout the city. Lin et al. (Lin et al., 2018) developed a sophisticated contextual multi-agent reinforcement learning framework that combines contextual deep Q-learning with a contextual multi-agent actor-critic approach. This framework facilitates explicit coordination between a large number of vehicles operating on a large-scale, on-demand ridesharing platform. By incorporating both learning techniques, the framework is able to enhance the decision-making processes of each agent (vehicle) while enabling them to work together seamlessly. In another work, Teng et al. (Tang et al., 2019) developed a cerebellar value network for optimizing multi-driver order dispatching in ridesharing systems. This work aims to improve the stability of state value iteration when employing nonlinear function approximators and demonstrate that this new approach substantially enhances driver income and passenger experience metrics. Al-Abbasi et al. (Al-Abbasi et al., 2019) developed a model-free method leveraging deep Q-network (DQN) techniques to determine optimal dispatch policies through interaction with the environment in a ridesharing system.
@@
Fairness in Ridesharing. Fairness is a critical factor in matching scenarios, particularly in multi-sided markets where preferences play a significant role (Garfinkel, 1971; Lorenz, 1905; Dickerson et al., 2014). In the context of ridesharing, recent literature has raised concerns about the fairness of the assignments. Brown et al. (Brown, 2018) shows the unfair treatment of riders by ridesharing companies, which leads to higher trip cancellation rates for a few groups of riders. Another study indicates that income inequality among ridesharing drivers may prevent them from earning a living wage (Graham and Joe, 2017). Another work aims to optimize long-term efficiency and fairness in ridesharing platforms through joint order dispatching and driver repositioning (Sun et al., 2024). Another recent work (Sahebdel et al., 2024) showed a trade-off between reducing carbon emissions and sacrificing the fairness of utilities among drivers. Similarly, another study (Shi et al., 2021b) demonstrated optimization in driver utilities and fairness without considering emissions. This study used reinforcement learning techniques to evaluate the expected achievable utilities in different city areas. Despite the success of this approach in ensuring equity among drivers’ utilities, it did not address eco-friendly ride assignments with low carbon emissions.
@@
Ensuring fair assignments in ridesharing systems is challenging due to drivers’ and riders’ inherent spatiotemporal dynamics. Drivers can earn varying amounts of utility based on their assignments, while riders may experience differing waiting times. These disparities can lead to perceptions of unfairness, complicating the assignment process. To tackle these challenges, our work introduces a new fairness metric for driver earnings that considers the spatiotemporal dynamics inherent in ridesharing systems. This metric is designed to ensure a more equitable distribution of earnings among drivers, considering the dynamic nature of the system. Building on this metric, we have developed a reinforcement learning-based approach assignment algorithm that reduces carbon emissions while providing a fair distribution of driver utilities compared to the state-of-the-art assignment algorithm (Sahebdel et al., 2024). The algorithm leverages historical data and dynamic learning to create effective and equitable assignment strategies, addressing immediate and long-term objectives in the ridesharing ecosystem.
3. Problem Formulation
@@
In ridesharing platforms, riders set the pickup and dropoff locations for a ride request; then, the platform matches the request to an available driver and estimates the trip duration and waiting time based on factors such as trip distance, traffic congestion, and the distance between the assigned driver and the rider’s pickup location. An available driver in the system refers to a driver who is either idle or soon to be available within the specified time to serve a request. The emission associated with a ride request $r$ is influenced by the unit emissions of the driver’s vehicle, the deadhead distance, the trip distance, and additional factors such as traffic congestion, road conditions, etc. For simplicity, we estimate the carbon emissions produced during the servicing of the ride request $r$ by driver $v$ using the equation below.
(1) $E_{v,r}=(d_{v,r}^{(D)}+d_{r}^{(T)})\cdot e_{v},$
@@
(2) $U_{v,r}=\sum_{r^{\prime}\in\mathcal{V}_{r}}(d_{r^{\prime}}^{(T)}-d_{v,r^{\prime}}^{(D)})\cdot x_{v,r^{\prime}},$
@@
$\text{Fairness}:=-\max_{v}(U_{v})+\min_{v}(U_{v}),$
(3e) s.t., $\sum_{v}x_{v,r}\in\{0,1\}\qquad\forall r,$
@@
(3f) vars., $\ \ \quad x_{v,r}\in\{0,1\}\qquad\forall v,r.$